# Study of mass outflows from magnetized accretion disks around rotating black holes with thermal conduction

### v1

@@ Abstract
We examine mass outflows from a relativistic, viscous, advective, and magnetized accretion disk around a rotating black hole in presence of thermal conduction. We consider the disk is primarily threaded by the toroidal component of the magnetic field and an effective potential satisfactorily mimicked the spacetime geometry around the rotating black hole. With this, we self-consistently solve the coupled governing equations for inflow and outflow and compute the mass outflow rate $R_{\dot{\rm}m}$ (ratio of mass flux of inflow to outflow) in terms of the inflow parameters, namely energy ($\mathcal{E}$), angular momentum ($\lambda$), plasma-$\beta$ and conduction parameter ($\Upsilon_{\rm s}$) around weakly rotating ($a_{\rm k}\rightarrow 0$) as well as rapidly rotating ($a_{\rm k}=0.99$) black holes. Our findings reveal that the present formalism admits coupled inflow-outflow solutions across a wide range of inflow parameters yielding substantial mass loss. We observe that $R_{\dot{\rm m}}$ monotonically increases with $\Upsilon_{\rm s}$, irrespective of black hole spin. We also find that for a fixed $\Upsilon_{\rm s}$, when energy, angular momentum, and magnetic field strength of the inflowing matter is increased, $R_{\dot{\rm m}}$ is enhanced resulting the outflows even more pronounced. We further estimate the maximum outflow rate ($R^{\rm max}_{\dot{\rm m}}$) by varying the inflow parameters and find that thermal conduction leads to maximum mass outflow rate $R^{\rm max}_{\dot{\rm m}}\sim 24\%$ for rapidly rotating black hole of spin $a_{\rm k}=0.99$. Finally, we employ our formalism to explain the kinetic jet power of $68$ radio-loud low-luminosity active galactic nuclei (LLAGNs), indicating that it is potentially promising to account for the observed jet power of substantial number of LLAGNs.
@@ I Introduction
Motivated by this, we extend our previous work Jana and Das (2024), and study the accretion-ejection mechanism in a relativistic, viscous, advective, magnetized accretion disk, incorporating the effects of thermal conduction. We consider accretion flow predominantly threaded by toroidal magnetic fields Oda et al. (2007) and adopt an effective potential to mimic the space-time around a rotating black holes Dihingia et al. (2018). We self-consistently solve the coupled inflow and outflow equations around the rotating black hole of spin $a_{\rm k}$ and quantify the mass outflow rate ($R_{\dot{\rm m}}$) in terms of inflow parameters, namely conduction parameter ($\Upsilon_{\rm s}$), energy ($\mathcal{E}$), angular momentum ($\lambda$), and plasma-$\beta$ (measure of magnetic fields strength). Our findings indicate that thermal conduction significantly influences $R_{\dot{\rm m}}$; as thermal conduction increases, so does the outflow rate. Furthermore, we calculate the maximum outflow rate ($R^{\rm max}_{\dot{\rm m}}$) by freely varying the model parameters and observe that $R^{\rm max}_{\dot{\rm m}}$ remains higher for a rapidly rotating black hole ($a_{\rm k}=0.99$) compared to a non-rotating black hole ($a_{\rm k}=0.0$). Finally, we discuss the implications of our findings in the context of jet kinetic power observed in low luminosity AGNs (LLAGNs).
@@ II.1 Governing equations for accretion
We begin with a relativistic, low angular momentum, viscous, magnetized, advective accretion flow around a rotating BH, incorporating thermal conduction as the heat transfer mechanism. To model the magnetic field structure, we refer to numerical simulations Hirose et al. (2006); Machida et al. (2006); Johansen and Levin (2008), which demonstrate that the magnetic fields inside the disk are turbulent in nature and primarily dominated by the toroidal component. Following these simulation works, the total magnetic fields can be decomposed into mean field, $B$ = (00, $\langle B_{\phi}\rangle$, 00) and fluctuating field, $\delta B$ = ($\delta B_{x}$, $\delta B_{\phi}$, $\delta B_{z}$). Here, $\langle$ $\rangle$’ corresponds to the azimuthal average. Upon azimuthal averaging, we assume that the fluctuating terms vanish ($\langle\delta B\rangle=0$). Therefore, the azimuthal component becomes dominant over the radial and vertical component, i.e, $|\langle B_{\phi}\rangle+\delta B_{\phi}|$ >>much-greater-than$>$ > $|\delta B_{x}|$ and $|\delta B_{z}|$. This ultimately leads to the azimuthally averaged magnetic field as $\langle B\rangle$ = $\hat{\phi}$ $\langle B_{\phi}\rangle$ Oda et al. (2007). Further, to account for general relativistic effect, we adopt the recently introduced pseudo-potential Dihingia et al. (2018), which satisfactorily replicates the space-time geometry around a rotating BH. The effective potential at the disk equatorial plane is given by,
@@
$v\frac{dv}{dx}+\frac{1}{h\rho}\frac{dP}{dx}+\frac{d\Psi_{\rm{e}}^{\rm{eff}}}{dx}+\frac{\langle B_{\phi}^{2}\rangle}{4\pi x\rho}=0,$ (2)
Here, $x$, $v$, $h$, and $\rho$ represent the radial distance, radial velocity, specific enthalpy, and mass density, respectively. The specific enthalpy is calculated as $h=(\epsilon+P_{\rm gas})/\rho$, where $\epsilon$ denotes the internal energy of the flow. The total isotropic pressure $P$ combines both gas pressure ($P_{\rm gas}$) and magnetic pressure ($P_{\rm mag}$), such that $P=P_{\rm gas}+P_{\rm mag}$. For simplicity, we consider that ions are predominantly protons, and the electron number density ($n_{\rm e}$) equals the ion number density ($n_{\rm p}$), leading to $n_{\rm e}=n_{\rm p}$. Subsequently, assuming a single temperature for both ions and electrons, we express $P_{\rm gas}=2\rho\Theta/\tau$, where $\tau=1+m_{\rm p}/m_{\rm e}$ and $\Theta$ is the dimensionless temperature given as $\Theta=k_{\rm B}T/{m_{\rm e}c^{2}}$. Here, $k_{\rm B}$ is the Boltzmann constant, $T$ is the temperature in Kelvin, and $m_{\rm e}$ and $m_{\rm p}$ are the electron and proton masses, respectively. Further, the magnetic pressure is obtained as $P_{\rm mag}=\frac{\langle B_{\phi}^{2}\rangle}{8\pi}$, where $\langle B_{\phi}^{2}\rangle$ denotes the azimuthal average of the square of toroidal component of the magnetic fields. We define the plasma-$\beta$ as the ratio of gas pressure to magnetic pressure, given by $\beta=P_{\rm gas}/P_{\rm mag}$, and estimate the magnetic strength as $\langle B_{\phi}^{2}\rangle=8\pi P_{\rm g}/\beta$.
@@
In equation (5), we consider that the vertically integrated total stress is dominated by the ${x\phi}$-component of the Maxwell stress ($T_{x\phi}$) over the other components. Following Machida et al. (2006), we calculate $T_{{\rm x}\phi}$ in presence of significant radial motion of the accreting matter as Chakrabarti (1996),
@@
where $W~{}(=2PH)$ is the vertically integrated total pressure Matsumoto et al. (1984), and $\alpha_{\rm B}$ (ratio of Maxwell stress to the total pressure) is the constant of proportionality commonly known as viscosity parameter.
@@
$\frac{v\Sigma}{\Gamma-1}\left(\frac{1}{\rho}\frac{dP_{\rm{g}}}{dx}-\frac{\Gamma P_{\rm{g}}}{\rho^{2}}\frac{d\rho}{dx}\right)=Q^{-}-Q^{+}-Q^{\rm cond},$ (7)
@@
$Q^{\rm{vis}}=\frac{\langle B_{x}B_{\phi}\rangle}{4\pi}xH\frac{d\Omega}{dx}=-\alpha_{\rm{B}}\left(W+\Sigma v^{2}\right)x\frac{d\Omega}{dx}.$ (8)
@@
where $F_{\rm s}$ represents the saturated conduction flux, defined as $F_{\rm s}=5\Upsilon_{\rm{s}}\rho\left(\frac{P_{\rm gas}}{\rho}\right)^{3/2}$. Here, $\Upsilon_{\rm{s}}$ is the dimensionless saturated conduction parameter (hereafter conduction parameter) that regulates the effect of thermal conduction within the disk and it takes values in the range $10\leq\Upsilon_{\rm s}<$ 1. Furthermore, since this study focuses on radiatively inefficient hot accretion flows characterized by low accretion rates, we neglect the cooling effects and set $Q^{-}=0$.
@@
where $f$ is defined in terms of the temperature as,
@@ II.2 Governing equations for outflows
The complex interaction between accretion and ejection near black holes provides insights into mass loss activities, and hence we assume that outflows originate from the accretion disk and move along the black hole rotation axis. As a port of the inflowing matter deflects as outflows, leading to mass loss from the system, these outflows remain tenuous in nature. Because of this, we disregard the differential rotation and neglect viscosity in the outflows. In addition, as the toroidal component of the magnetic field is assumed to be dominant and the matter is being ejected vertically, we neglect the magnetic fields within the outflows for simplicity. Furthermore, we adopt the polytropic equation of state, $P_{\rm j}=K_{\rm j}\rho^{\Gamma}_{\rm j}$, where the subscript ‘$\rm j$’ represents jet variables, $K_{\rm j}$ measures the entropy of the outflowing matter and $\Gamma$ is the adiabatic index which is determined based on temperature using the REoS (equation 12, 13). Based on these considerations, the governing equations for outflows are as follows:
@@
In equation (18), $\mathcal{E}_{\rm j}$, $v_{\rm j}$, and $h_{\rm j}$ denote energy, velocity, and enthalpy of the outflows. Here, $\Psi^{\rm eff}$ refers the effective potential and is given by Dihingia et al. (2018) ,
@@
where $\mathcal{Y}=(r^{2}_{\rm j}+a^{2}_{\rm k})^{2}-\Delta a_{\rm k}^{2}(x_{\rm j}/r_{\rm j})^{2}$, $\mathcal{Z}={r^{2}_{\rm j}+a^{2}_{\rm k}\Big{(}1-(x_{\rm j}/r_{\rm j})^{2}\Big{)}}$ and $r_{\rm j}~{}\left(=\sqrt{x^{2}_{\rm j}+z^{2}_{\rm j}}\right)$ is the spherical radius of outflow Jana and Das (2024). In equation (19), $\dot{M}_{\rm out}$ is the outflowing mass rate and $\mathcal{A}_{\rm j}$ is the area function. In this work, we adopt a jet geometry where the outflowing matter is confined between two boundaries, namely centrifugal wall (CB) and funnel wall (FW) Molteni et al. (1996); Das and Chakrabarti (2008); Aktar et al. (2015); Jana and Das (2024). The CB is defined as $\left(\frac{d\Psi^{\rm eff}}{dx}\right)_{r_{\rm CB}}=0$ and the radius of funnel wall is determined using $\Psi^{\rm eff}|_{r_{\rm FW}}$ = 00. Subsequently, the jet coordinates are defined as, $x_{\rm j}=(x_{\rm CB}+x_{\rm FW})/2$ and $z_{\rm j}=z_{\rm CB}=z_{FW}$. With this, we calculate the area function as $\mathcal{A}=2\pi(x^{2}_{\rm CB}-x^{2}_{\rm FW})/\sqrt{1+(dx_{\rm j}/dz_{\rm j})^{2}}$, where $\sqrt{1+(dx_{\rm j}/dz_{\rm j})^{2}}$ represents the projection factor Kumar and Chattopadhyay (2013). Similar to accretion, we carry out the critical point analysis and obtain the outflow velocity ($v_{\rm jc}$) at the outflow critical point ($r_{\rm jc}$) from the critical point conditions as Das and Chattopadhyay (2008); Jana and Das (2024),
@@
where the subscript ‘$\rm jc$’ indicates the jet quantities at $r_{\rm jc}$ along with $r=r_{\rm CB}$. Using equation (21), the outflow temperature at $r_{\rm jc}$ is determined by numerically solving the following equation:
$\sqrt{\frac{\Theta_{\rm jc}}{Nh_{\rm jc}}\Big{(}\frac{dh_{\rm j}}{d\Theta_{\rm j}}\Big{)}_{\Theta_{\rm jc}}}=v_{\rm jc}.$ (22)
@@ II.3 Disk-jet connection
During the accretion process, matter experiences two opposing forces, namely the gravitational pull and centrifugal repulsion. When these forces become comparable at the vicinity of the black hole, the supersonic inflow decelerates causing matter to accumulate. However, this accumulation does not continue indefinitely, instead, once a critical threshold is reached, centrifugal repulsion triggers a shock transition Frank et al. (2002). The resulting shock compression makes the post-shock flow — also known as the post-shock corona (PSC) — dense and hot, forming a puffed-up, torus-like effective boundary around the black hole. The excess thermal gradient force across the shock front then drives a fraction of the infalling matter vertically outward as bipolar outflows. Since this ejection mechanism is linked to the accretion process through shock formation, we apply shock conditions to solve the inflow and outflow equations. In presence of mass loss, the shock conditions Landau and Lifshitz (1959); Das and Chattopadhyay (2008); Sarkar et al. (2018) are as follows: (a) the mass conservation $\dot{M}_{+}=\dot{M}_{-}-\dot{M}_{\rm out}=\dot{M}_{-}(1-R_{\dot{\rm m}})$, (b) the energy conservation: $\mathcal{E}_{+}=\mathcal{E}_{-}$, (c) the momentum conservation: $W_{+}+\Sigma_{+}v^{2}_{+}=W_{-}+\Sigma_{-}v^{2}_{-}$, and (d) the magnetic flux conservation $\dot{\Phi}_{+}=\dot{\Phi}_{-}$. Here, $R_{\dot{\rm m}}~{}(=\dot{M}_{\rm out}/\dot{M}_{-})$ is the outflow rate, and the symbols ‘$+$’ and ‘--’ refer the quantities immediately after and before the shock transition.

### v2

@@ Abstract
We examine mass outflows from a low-angular momentum, viscous, advective, and magnetized accretion disk around a rotating black hole in presence of thermal conduction. We consider the disk is primarily threaded by the toroidal component of the magnetic field and an effective potential satisfactorily mimicked the spacetime geometry around the rotating black hole. With this, we self-consistently solve the coupled governing equations for inflow and outflow and compute the mass outflow rate $R_{\dot{m}}$ (ratio of mass flux of inflow to outflow) in terms of the inflow parameters, namely energy ($\mathcal{E}$), angular momentum ($\lambda$), plasma-$\beta$ and conduction parameter ($\Upsilon_{\rm s}$) around weakly rotating ($a_{\rm k}\rightarrow 0$) as well as rapidly rotating ($a_{\rm k}=0.99$) black holes. Our findings reveal that the present formalism admits coupled inflow-outflow solutions across a wide range of inflow parameters yielding substantial mass loss. We observe that $R_{\dot{m}}$ monotonically increases with $\Upsilon_{\rm s}$, irrespective of black hole spin. We also find that for a fixed $\Upsilon_{\rm s}$, when energy, angular momentum, and magnetic field strength of the inflowing matter is increased, $R_{\dot{\rm m}}$ is enhanced resulting the outflows even more pronounced. We further estimate the maximum outflow rate ($R^{\rm max}_{\dot{\rm m}}$) by varying the inflow parameters and find that thermal conduction leads to maximum mass outflow rate $R^{\rm max}_{\dot{\rm m}}\sim 25\%$ for rapidly rotating black hole of spin $a_{\rm k}=0.99$. Finally, we employ our formalism to explain the kinetic jet power of $68$ radio-loud low-luminosity active galactic nuclei (LLAGNs), indicating that it is potentially promising to account for the observed jet power of substantial number of LLAGNs.
@@ I Introduction
Motivated by this, we extend our previous work Jana and Das (2024), and study the accretion-ejection mechanism in a low angular momentum, viscous, advective, magnetized accretion disk, incorporating the effects of thermal conduction. We consider accretion flow predominantly threaded by toroidal magnetic fields Oda et al. (2007) and adopt an effective potential to mimic the space-time around a rotating black holes Dihingia et al. (2018). We self-consistently solve the coupled inflow and outflow equations around the rotating black hole of spin $a_{\rm k}$ and quantify the mass outflow rate ($R_{\dot{\rm m}}$) in terms of inflow parameters, namely conduction parameter ($\Upsilon_{\rm s}$), energy ($\mathcal{E}$), angular momentum ($\lambda$), and plasma-$\beta$ (measure of magnetic fields strength). Our findings indicate that thermal conduction significantly influences $R_{\dot{\rm m}}$; as thermal conduction increases, so does the outflow rate. Furthermore, we calculate the maximum outflow rate ($R^{\rm max}_{\dot{\rm m}}$) by freely varying the model parameters and observe that $R^{\rm max}_{\dot{\rm m}}$ remains higher for a rapidly rotating black hole ($a_{\rm k}=0.99$) compared to a non-rotating black hole ($a_{\rm k}=0.0$). Finally, we discuss the implications of our findings in the context of jet kinetic power observed in low luminosity AGNs (LLAGNs).
@@ II.1 Governing equations for accretion
We begin with a low angular momentum, viscous, magnetized, advective accretion flow around a rotating BH, incorporating thermal conduction as the heat transfer mechanism. To model the magnetic field structure, we refer to numerical simulations Hirose et al. (2006); Machida et al. (2006); Johansen and Levin (2008), which demonstrate that the magnetic fields inside the disk are turbulent in nature and primarily dominated by the toroidal component. Following these simulation works, the total magnetic fields can be decomposed into mean field, $B$ = (0, $\langle B_{\phi}\rangle$, 0) and fluctuating field, $\delta B$ = ($\delta B_{x}$, $\delta B_{\phi}$, $\delta B_{z}$). Here, $\langle$ $\rangle$’ corresponds to the azimuthal average. Upon azimuthal averaging, we assume that the fluctuating terms vanish ($\langle\delta B\rangle=0$). Therefore, the azimuthal component becomes dominant over the radial and vertical component, i.e, $|\langle B_{\phi}\rangle+\delta B_{\phi}|$ >>>> $|\delta B_{x}|$ and $|\delta B_{z}|$. This ultimately leads to the azimuthally averaged magnetic field as $\langle B\rangle$ = $\hat{\phi}$ $\langle B_{\phi}\rangle$ Oda et al. (2007). Further, to account for general relativistic effect, we adopt the recently introduced pseudo-potential Dihingia et al. (2018), which satisfactorily replicates the space-time geometry around a rotating BH. The effective potential at the disk equatorial plane is given by,
@@
$v\frac{dv}{dx}+\frac{1}{\rho}\frac{dP}{dx}+\frac{d\Psi_{\rm{e}}^{\rm{eff}}}{dx}+\frac{\langle B_{\phi}^{2}\rangle}{4\pi x\rho}=0,$ (2)
Here, $x$, $v$, and $\rho$ represent the radial distance, radial velocity, and mass density, respectively. The total isotropic pressure $P$ combines both gas pressure ($P_{\rm gas}$) and magnetic pressure ($P_{\rm mag}$), such that $P=P_{\rm gas}+P_{\rm mag}$. For simplicity, we consider that ions are predominantly protons, and the electron number density ($n_{\rm e}$) equals the ion number density ($n_{\rm p}$), leading to $n_{\rm e}=n_{\rm p}$. Subsequently, assuming a single temperature for both ions and electrons, we express $P_{\rm gas}=2\rho\Theta/\tau$, where $\tau=1+m_{\rm p}/m_{\rm e}$ and $\Theta$ is the dimensionless temperature given as $\Theta=k_{\rm B}T/{m_{\rm e}c^{2}}$. Here, $k_{\rm B}$ is the Boltzmann constant, $T$ is the temperature in Kelvin, and $m_{\rm e}$ and $m_{\rm p}$ are the electron and proton masses, respectively. Further, the magnetic pressure is obtained as $P_{\rm mag}=\frac{\langle B_{\phi}^{2}\rangle}{8\pi}$, where $\langle B_{\phi}^{2}\rangle$ denotes the azimuthal average of the square of toroidal component of the magnetic fields. We define the plasma-$\beta$ as the ratio of gas pressure to magnetic pressure, given by $\beta=P_{\rm gas}/P_{\rm mag}$, and estimate the magnetic strength as $\langle B_{\phi}^{2}\rangle=8\pi P_{\rm gas}/\beta$.
@@
In equation (5), we consider that the vertically integrated total stress is dominated by the $x\phi$}-component of the Maxwell stress ($T_{x\phi}$) over the other components. Following Machida et al. (2006), we calculate $T_{x\phi}$ in presence of significant radial motion of the accreting matter as Chakrabarti (1996),
@@
where W(=2PH)W~(=2PH) is the vertically integrated total pressure Matsumoto et al. (1984), and $\alpha_{\rm B}$ (ratio of Maxwell stress to the total pressure) is the constant of proportionality commonly known as viscosity parameter.
@@
$\frac{v\Sigma}{\Gamma-1}\left(\frac{1}{\rho}\frac{dP_{\rm gas}}{dx}-\frac{\Gamma P_{\rm gas}}{\rho^{2}}\frac{d\rho}{dx}\right)=Q^{-}-Q^{+}-Q^{\rm cond},$ (7)
@@
$Q^{\rm+}=\frac{\langle B_{x}B_{\phi}\rangle}{4\pi}xH\frac{d\Omega}{dx}=-\alpha_{\rm{B}}\left(W+\Sigma v^{2}\right)x\frac{d\Omega}{dx}.$ (8)
@@
where $F_{\rm s}$ represents the saturated conduction flux, defined as $F_{\rm s}=5\Upsilon_{\rm{s}}\rho\left(\frac{P_{\rm gas}}{\rho}\right)^{3/2}$. Here, $\Upsilon_{\rm{s}}$ is the dimensionless saturated conduction parameter (hereafter conduction parameter) that regulates the effect of thermal conduction within the disk and it takes values in the range $0\leq\Upsilon_{\rm s}<1$. Furthermore, since this study focuses on radiatively inefficient hot accretion flows characterized by low accretion rates, we neglect the cooling effects and set $Q^{-}=0$.
@@
where $\epsilon$ denotes the internal energy of the flow and $f$ is defined in terms of the temperature as,
@@ II.2 Governing equations for outflows
The complex interaction between accretion and ejection near black holes provides insights into mass loss activities, and hence we assume that outflows originate from the accretion disk and move along the black hole rotation axis. As a part of the inflowing matter deflects as outflows, leading to mass loss from the system, these outflows remain tenuous in nature. Because of this, we disregard the differential rotation and neglect viscosity in the outflows. In addition, as the toroidal component of the magnetic field is assumed to be dominant and the matter is being ejected vertically, we neglect the magnetic fields within the outflows for simplicity. Furthermore, we adopt the polytropic equation of state, $P_{\rm j}=K_{\rm j}\rho^{\Gamma}_{\rm j}$, where the subscript $\rm j$’ represents jet variables, $K_{\rm j}$ measures the entropy of the outflowing matter and $\Gamma$ is the adiabatic index which is determined based on temperature using the REoS (equation 12, 13). Based on these considerations, the governing equations for outflows are as follows:
@@
In equation (18), $\mathcal{E}_{\rm j}$, $v_{\rm j}$ denote energy and velocity of the outflows, and $C_{\rm sj}~(=\sqrt{\Gamma P_{\rm j}/\rho_{\rm j}})$ is the sound speed of the outflowing matter. Here, $\Psi^{\rm eff}$ refers the effective potential and is given by Dihingia et al. (2018) ,
@@
where $\mathcal{Y}=(r^{2}_{\rm j}+a^{2}_{\rm k})^{2}-\Delta a_{\rm k}^{2}(x_{\rm j}/r_{\rm j})^{2}$, $\mathcal{Z}={r^{2}_{\rm j}+a^{2}_{\rm k}\Big{(}1-(x_{\rm j}/r_{\rm j})^{2}\Big{)}}$ and $r_{\rm j}~\left(=\sqrt{x^{2}_{\rm j}+z^{2}_{\rm j}}\right)$ is the spherical radius of outflow Jana and Das (2024). In equation (19), $\dot{M}_{\rm out}$ is the outflowing mass rate and $\mathcal{A}_{\rm j}$ is the area function. In this work, we adopt a jet geometry where the outflowing matter is confined between two boundaries, namely centrifugal wall (CB) and funnel wall (FW) Molteni et al. (1996); Das and Chakrabarti (2008); Aktar et al. (2015); Jana and Das (2024). The CB is defined as $\left(\frac{d\Psi^{\rm eff}}{dx}\right)_{r_{\rm CB}}=0$ and the radius of funnel wall is determined using $\Psi^{\rm eff}|_{r_{\rm FW}}$ = 0. Subsequently, the jet coordinates are defined as, $x_{\rm j}=(x_{\rm CB}+x_{\rm FW})/2$ and $z_{\rm j}=z_{\rm CB}=z_{FW}$. With this, we calculate the area function as $\mathcal{A}=2\pi(x^{2}_{\rm CB}-x^{2}_{\rm FW})/\sqrt{1+(dx_{\rm j}/dz_{\rm j})^{2}}$, where $\sqrt{1+(dx_{\rm j}/dz_{\rm j})^{2}}$ represents the projection factor Kumar and Chattopadhyay (2013). Similar to accretion, we carry out the critical point analysis and obtain the outflow velocity ($v_{\rm jc}$) at the outflow critical point ($r_{\rm jc}$) from the critical point conditions as Das and Chattopadhyay (2008); Jana and Das (2024),
@@
where the subscript $\rm jc$’ indicates the jet quantities at $r_{\rm jc}$ along with $r=r_{\rm CB}$. Using equation (21), the outflow temperature at $r_{\rm jc}$ is determined by numerically solving the following equation:
$\sqrt{\frac{\Theta_{\rm jc}}{N}\Big{[}\frac{d}{d\Theta_{\rm j}}\left(NC^{2}_{\rm sj}\right)\Big{]}_{\Theta_{\rm jc}}}=v_{\rm jc}.$ (22)
@@ II.3 Disk-jet connection
During the accretion process, matter experiences two opposing forces, namely the gravitational pull and centrifugal repulsion. When these forces become comparable at the vicinity of the black hole, the supersonic inflow decelerates causing matter to accumulate. However, this accumulation does not continue indefinitely, instead, once a critical threshold is reached, centrifugal repulsion triggers a shock transition Frank et al. (2002). The resulting shock compression makes the post-shock flow — also known as the post-shock corona (PSC) — dense and hot, forming a puffed-up, torus-like effective boundary around the black hole. The excess thermal gradient force across the shock front then drives a fraction of the infalling matter vertically outward as bipolar outflows. Since this ejection mechanism is linked to the accretion process through shock formation, we apply shock conditions to solve the inflow and outflow equations. In presence of mass loss, the shock conditions Landau and Lifshitz (1959); Das and Chattopadhyay (2008); Sarkar et al. (2018) are as follows: (a) the mass conservation $\dot{M}_{+}=\dot{M}_{-}-\dot{M}_{\rm out}=\dot{M}_{-}(1-R_{\dot{\rm m}})$, (b) the energy conservation: $\mathcal{E}_{+}=\mathcal{E}_{-}$, (c) the momentum conservation: $W_{+}+\Sigma_{+}v^{2}_{+}=W_{-}+\Sigma_{-}v^{2}_{-}$, and (d) the magnetic flux conservation $\dot{\Phi}_{+}=\dot{\Phi}_{-}$. Here, $R_{\dot{\rm m}}~(=\dot{M}_{\rm out}/\dot{M}_{-})$ is the outflow rate, and the symbols ‘++’ and ‘-’ refer the quantities immediately after and before the shock transition.